Grade 11Mechanics

Two solid spheres of radius 10cm and masses 800kg and 600kg are at a distance 0.25m apart.Find gravitational field intensity at a point at a distance 0.2m from the 800 kg sphere and 0.15m from the 600kg sphere and is not collinear.?

Deva Vikaas
8 Years agoGrade 11

1 Answer

Sweety Negi
7 Years ago
Let Ebe the gravitational field intensity at point P due to 800 kg sphere at A.
Then, EA = G 800/(0.2)= 2 x 10along PA
Let EB be the gravitational field intensity at P due yo 600 kg sphere at B
Then, EB = G 600/(0.15)= 80000G/3 along PB
 
The angle between EA and EB is 90°. [•.•  AB= AP+ BP]
If E be the magnitude of the resultant intensity, then
E = √(EA+ EB2)
   = √(2 x 10G)+ (80000G/3)2
   =G √4 x 10+ 64/9 x 108
    = 6.67 x 10-11 x 2 x 10√1 + 16/9
   = 6.67 x 2 x 5/3 x 10-7 N kg-1
E = 2.22 x 10-6 N kg-1